A Styrofoam cylinder, filled with water, sits on a table. You then poke a small hole 

through the side of the cylinder, 20 cm below the top of the water surface. What is 

the speed of the fluid emerging from the hole?


1
Expert's answer
2021-03-29T08:58:33-0400

Let's write the Bernoulli equation (indexes 1 and 2 corresponds to the top and bottom, respectively):


"\dfrac{\rho v_1^2}{2}+\rho gh_1+p_1=\dfrac{\rho v_2^2}{2}+\rho gh_2+p_2."

Since "v_1=0""h_2=0" and "p_1=p_2", we get:


"\rho gh_1=\dfrac{\rho v_2^2}{2},""v_2=\sqrt{2gh_1}=\sqrt{2\cdot9.8\ \dfrac{m}{s^2}\cdot0.2\ m}=2.0\ \dfrac{m}{s}."